OpenStudy (anonymous):

y^3-11y^2+28y factor the trinomial completely

OpenStudy (anonymous):

y^3-11y^2+28y =y(y^2-11y+28) =y(y^2-7y-4y+28) =y[y(y-7)-4(y-7)] =y(y-7)(y-4)